OpenStudy (anonymous):

Find an explicit formula and a recurrence formula for the following sequence. {1, -1/2, 3, -1/4, 5...}. I've been at it for awhile and just can't seem to come up with the correct formulas.

OpenStudy (kainui):

Yeah this is tricky looking. What sorta patterns have you found and tried out? I can help show you how to adjust them to work I think that'd be easier for me to explain and better for you too

OpenStudy (anonymous):

so far I've only found complicated ones that haven't worked, and I've erased them (which I realize now was not a good idea) but the last one I have is \[\frac{n}{-n} * \frac{1}{n} * (-(n)^{2})\]

OpenStudy (anonymous):

Are there any strategies to coming up with the formulas for these?

OpenStudy (kainui):

Yeah it's sorta hard to figure out a strategy for guessing these, but I like to kind of lay out what I know as best as I can. It might help to list stuff. So when I see 1, -1/2, 3, -1/4, 5... I see, ok it's kinda like 1, 2, 3, 4, 5, ... And that much I know I can write as n. So maybe I can try to alter this to get it to work. It looks like there are two other things going on here, we're getting this thing to flip flop between n and 1/n AND it's also flip flopping between + and -. So I'd try to focus on this part next. I think + vs - flip-flopping is easier personally so let's try to figure out how to make this sequence: 1, -1, 1, -1, 1, ... Then when we find it, we can figure out how to attach it the sequence 1,2,3,4,5,...

OpenStudy (anonymous):

Yeah, I was actually just thinking about that as well. So a -(1/n) would almost work, except it would give me the inverse of each index. I feel like that is a good start though?

OpenStudy (kainui):

Yeah definitely!

ganeshie8 (ganeshie8):

starting the index from \(n=1\), \(a_{2n+1} = 2n+1\) \(a_{2n} = -\dfrac{1}{2n} \) combine them and get \(a_n = (1+(-1)^n)(\dfrac{-1}{n}) + (1-(-1)^n)\dfrac{2n+1}{2} \)

OpenStudy (kainui):

Also, are you familiar with writing fractions as exponents like this? \[\frac{1}{n} = n^{-1}\] I think this will make the alternating part easier for you to figure out since you probably know that a number itself is raised to the power 1: \[n=n^1\] So I'll just leave these here to give you some ideas to think about, try to figure it out

OpenStudy (anonymous):

Yes, I'm familiar with that way of writing fractions, it definitely makes things like integrating a lot easier. Thanks, I'll try it out

ganeshie8 (ganeshie8):

i think the key thing here in writing an explicit formula for interleaved sequences is in noting below little trick: \( 1+(-1)^n\) is \(0\) when \(n\) is odd and it is \(2\) when \(n\) is even

OpenStudy (anonymous):

I'm trying to see how that works in the previous equation you posted above. Wouldn't that make it equal to zero any time that exponent is odd?

ganeshie8 (ganeshie8):

First of all, do you see why breaking the given sequence into two subsequences is easy here ?

ganeshie8 (ganeshie8):

Consider below two subsequences : 1) sequence formed by taking only odd indexed terms 2) sequence formed by taking only even indexed terms

OpenStudy (anonymous):

Yes, that part made sense to me

ganeshie8 (ganeshie8):

Is it easy to write explicit formulas for each of those sequences indpendently ?

ganeshie8 (ganeshie8):

what are the explicit formulas ?

OpenStudy (anonymous):

I think \[a _{2n+1} = n\] and \[a _{2n} = \frac{-1}{n}\]

ganeshie8 (ganeshie8):

looks we have a small problem with indexing other than that you look good!

ganeshie8 (ganeshie8):

let me rephrase like this : when \(n=2k-1\) is odd, the term is same as the index itself : \(n\) when \(n=2k\) is even, the term is \(-\dfrac{1}{n}\)

ganeshie8 (ganeshie8):

see if that looks okay

OpenStudy (anonymous):

Oh, ok. Yes I think I see how for the odd mind was 2k+1 when it should be 2k-1 right?

ganeshie8 (ganeshie8):

Yes, next see how you can make use of the \(\dfrac{1+(-1)^n}{2}\) thingy

ganeshie8 (ganeshie8):

Notice again, above expression is \(0\) when \(n\) is odd, and it is \(1\) when \(n\) is even

OpenStudy (anonymous):

So it would be 1/2 when n is odd and 1 when n is even, right?

ganeshie8 (ganeshie8):

Nope, whats the value of (-1)^n when n is odd ?

ganeshie8 (ganeshie8):

whats the value of 1 + (-1)^n when n is odd ?

OpenStudy (anonymous):

Oh, my bad, I see it's zero when n is odd